OpenStudy (anonymous):

if f(x)= 3x^2-bx+4 and f(-1)=12 what is the value of b? [show steps ]

OpenStudy (mimi_x3):

3(-1)^2 -(-1)b + 4 = 12 3+b = 12 b = 12-3 b = 9

OpenStudy (lalaly):

\[f(-1)=3(-1)^2+b(-1)+4\]\[f(-1)=3+b+4 =12\]\[3+b+4=12 \rightarrow 7+b=12 \rightarrow b=5\]

OpenStudy (lalaly):

in the first step i did, 3(-1)^2 -b(-1) not +b(-1)

OpenStudy (mimi_x3):

woops , I made a stupid mistake xD . I forgot the 4 argh